Assuming the earth is a sphere whose radius is 6400 kilometer.Find the distance along a 3(degree)arc at the equator of the earth's surface.
Answer by gonzo(654)

the plane of the equator is a circle with a radius of 6400*km.
circumference of this circle would be 2*pi*r = 2*pi*6400km = 12800*pi.
since the circumference represents an arc of 360 degrees, a 3 degree arc would be 3/360 of this.
3/360 * 12800*pi = 106.666666667 * pi = 335.103*km rounded to the nearest thousandth of a km.